two angles are complimentary. the second angle is 62 degrees less than the first angle. what is the first angle?
Answer by checkley77(12569)

x+(x-62)=90
x+x-62=90
2x=90+62
2x=152
x=152/2
x=76 for the larger angle.
90-76=14 for the smaller angle.
Proof
76+(76-62)=90
76+14=90
90=90